# FlexLLM: A System for Co-Serving Large Language Model Inference and Parameter-Efficient Finetuning

### v1

@@ 6.2 Hybrid Token Scheduler
FlexLLM’s hybrid token scheduler schedules the token-level GPU computations in two steps. First, FlexLLM decides the schedule of available inference requests based on specific scheduling policies. FlexLLM uses Orca’s iteration-level scheduling [70] as the default scheduling policy, which maintains a fixed maximum batch size $b$ and dynamically replaces each completed request with a new one whenever available. Other scheduling policies, such as assigning different priorities to requests of different lengths, or partitioning the pre-filling phase of long input prompts into multiple pieces, are also feasible for FlexLLM. Second, after determining the inference schedule in each iteration, FlexLLM appends to the batch as many finetuning tokens as possible (i.e., sliding window size of $s$) for maximum resource utilization. The number of finetuning tokens that can be added is determined automatically using the formula $s=\operatornamewithlimits{arg\,max}f(c,s)\leq$SLO, where $f(\cdot,\cdot)$ is the latency estimation function and $c$ is the number of inference tokens that are running in the current iteration. The inference latency can be cheaply estimated by profiling the LLM execution costs in advance [46] so that the scheduling overhead is negligible compared to the per iteration latency.
Figure 11 illustrates the execution timeline of FlexLLM and shows an example of a token scheduling plan within a finetuning mini-batch $r_{A}$ under a particular inference requests arrival pattern (i.e., $r_{0}$ to $r_{9}$). To maintain the same finetuning semantic, FlexLLM’s scheduling follows the execution dependency between each mini-batch’s forward and backward pass. Within the forward pass, FlexLLM uses fused kernels to process both inference and finetuning tokens, avoiding additional kernel launch overhead. This is possible since the requests follow the same token-wise compute logic. Regarding the attention operator, finetuning requests follow the same logic and causal masking as inference requests in the pre-filling phase. For the backward pass, FlexLLM adapts a space-sharing approach, using two separate streams to handle inference and finetuning tokens.
@@ 7 Implementation
FlexLLM is built on top of FlexFlow Serve [44] with 9K lines of C++ and CUDA code. It comes with both C++ and Python bindings, and its Python interface is fully compatible with HuggingFace’s PEFT [41] and transformers [65] libraries. This means that users can easily serve models imported directly from HuggingFace, and even upload finetuned PEFTs to the HuggingFace Hub. Moreover, FlexLLM supports any PEFT models with LLAMA [62], GPT [9], OPT [74], Falcon [7], or MPT [61] backbones. In addition, we provide instructions on how to add support for custom models from HuggingFace. All the code was integrated into FlexFlow, available at https://github.com/flexflow/FlexFlow/. Further, FlexLLM is available to the users as a part of the FlexFlow pip package.
@@
To support token-level finetuning, it is necessary to accumulate key-value gradients for every preceding token in a given sequence during backpropagation. To achieve this, FlexLLM utilizes static allocation to reserve space for key-value gradient accumulation. During backpropagation, the gradients for keys and values of a specific segment are obtained, and the position of this segment in the overall sequence is determined. FlexLLM then accumulates the new key-value gradients of all subsequently scheduled segments. When a segment is scheduled during backpropagation, it includes the accumulated gradients brought by all future tokens.
Operator implementation
We have optimized kernels for NVIDIA GPUs using cuBLAS libraries for linear algebra operators and cuDNN for the other supported operators to run the most popular LLMs. Additionally, we have custom kernels for the remaining computations (such as layer normalization and SiLU activations). To reduce kernel launch overhead, we aggressively fuse customized kernels within and across operators. For example, we use a single kernel to compute the attention score for the inference tokens in the generative phase. We also fuse each residual summation with the following LayerNorm/RMSNorm operator and the Sigmoid, SiLU, and element-wise tensor multiplications from the LLaMA transformer layer.
@@ 8 Evaluation
We assessed the performance of FlexLLM by comparing it with existing approaches using three publicly available LLMs: OPT-13B [74], LLaMA-30B [62], and LLaMA-2-70B [63]. We obtained the pre-trained model parameters for the LLMs directly from their HuggingFace repositories [1]. Since existing LLM serving systems only support base LLMs and LoRA, our evaluation focuses on these two types of models to conduct a fair comparison. Like prior works, we applied LoRA (with a rank of 8) to both linear layers of each multi-layer perceptron block in each LLM. With this setup, we obtained, respectively, 15.6M, 30.5M, and 50.1M trainable parameters for the three LLMs.
@@ 8.1 End-to-end comparison
We first conducted an experiment to compare FlexLLM’s co-serving approach with existing approaches that utilize separate clusters to perform inference and finetuning. The baseline system we used employed HuggingFace (HF) PEFT [41] for finetuning and S-LoRA [57] for serving PEFT-based inference requests. In this experiment, we varied the inference workload and measured the finetuning throughput that different systems can achieve. We allocated a total of 4, 8, and 16 A100 GPUs for jointly serving inference and finetuning workloads for OPT-13B, LLaMA-30B, and LLaMA-2-70B, respectively. Each inference request had an average input prompt length of 256 and generated 64 tokens. We used a per-token inference latency of 50ms, 80ms, and 110ms as the SLA for inference requests. Each finetuning request had a sequence length of 1024 and used the Adam optimizer [31]. Figure 12 shows the result. We measured the number of inference tokens generated per second to quantify the inference workload. The baseline approach used a minimum number of GPUs to serve inference requests and allocated all remaining GPUs for finetuning PEFT models for a given inference workload.
Under lightweight inference workloads, FlexLLM achieves on-par performance with the baseline approach that uses separate clusters. This is because the baseline approach reserves most of the GPUs for fine-tuning PEFT models, using a very large fine-tuning batch size (e.g., 1024 for fine-tuning OPT-13B) when the inference workload is moderate. As a comparison, FlexLLM co-serves inference and finetuning using all GPUs and restricts the maximum finetuning batch size to satisfy the inference SLO. For example, to maintain a per-iteration co-serving latency of 50ms for OPT-13B, FlexLLM can process at most 384 tokens within an iteration.
Under medium inference workloads, FlexLLM achieves up to 1.6$\times$ higher finetuning throughput than the baseline. This performance improvement is achieved by utilizing more GPUs for finetuning and using available GPU cycles to process finetuning tokens more efficiently.
Finally, under heavy inference workloads, the existing approach has to allocate all GPUs for inference to cope with the high request arrival rates. As a result, it cannot make any progress on finetuning requests. In contrast, FlexLLM’s finetuning throughput only moderately decreases as we increase the inference workload and never falls below 80% of the peak finetuning throughput. Even when the existing approaches cannot make any progress, FlexLLM manages to perform well.
@@ 8.2 Memory Optimization
Figure 13: Detailed comparison of finetuning memory overheads.
Figure 14: Ablation study of memory saving techniques in FlexLLM.
Figure 13: Detailed comparison of finetuning memory overheads.
Figure 14: Ablation study of memory saving techniques in FlexLLM.
Figure 15: Comparing different scheduling strategies over FlexLLM.
Figure 15 illustrates the (finetuning-only) memory overhead analysis of FlexLLM across various PEFT and full finetuning methods on LLaMA-2-70B, with a sequence length 1024. We have compared the memory requirements of FlexLLM with HuggingFace PEFT and also included the memory usage for inference-only cases as a point of reference. In full-model finetuning, FlexLLM has the same memory requirement as HuggingFace. However, in the case of PEFT (LoRA, Adapter, and IA3), FlexLLM exhibits an impressive reduction of approximately $32\%$-$36\%$ in memory overhead compared to HuggingFace PEFT. This reduction brings FlexLLM closer to the memory overhead observed during inference, which is the lower boundary for finetuning procedures.
FlexLLM introduces a series of memory optimizations to reduce the memory overhead of PEFT finetuning and allow co-serving finetuning with inference workloads. To understand the effectiveness of FlexLLM memory optimizations, we perform an ablation study that incrementally turns off these optimizations (graph pruning, rematerialization, and token-level finetuning) and measures the memory requirements. Since all optimizations preserve model weights and reduce activation memory requirements, we only consider activation memory overheads in these experiments. Figure 15 shows the results. We analyze the activation memory requirements of FlexLLM for different finetuning methods on LLaMA-2-70B and a sequence length of 1024. FlexLLM saves $85\%$-$87\%$ memory requirements for activations compared to HF PEFT. The major improvement comes from graph pruning. With graph pruning only, FlexLLM achieves $71\%$-$74\%$ activation memory overhead reduction compared to HuggingFace PEFT. Rematerialization and token-level finetuning further reduce memory overhead by 00-$8\%$ and $4\%$-$10\%$, respectively.
8.3 GPU Scheduling
This section compares FlexLLM’s co-serving mechanism with two commonly used GPU scheduling strategies: temporal sharing and spatial sharing. These two baselines are implemented over FlexLLM but replace the co-serving mechanism with the corresponding approach. To conduct a fair comparison, we have enabled all memory optimizations for all scheduling strategies, so the difference in performance only comes from GPU scheduling. Temporal sharing schedules inference and finetuning computation in an interleaved fashion, as shown in Figure 3. Spatial sharing, conversely, uses separate CUDA streams to launch inference and finetuning computation in parallel. Figure 15 shows that even under a very high inference request arrival rate (i.e., 10 requests per second), the finetuning throughput of our co-serving approach only drops by around 10%, which is still far better than the temporal and spatial sharing strategies.
Refer to caption Refer to caption
Figure 16: Case study of FlexLLM’s system throughput and compute utilization for fluctuating inference workload.
8.4 Case Study
This section evaluates FlexLLM using a dynamic inference workload generated from a real online production trace at Microsoft Azure [55]. To ensure compatibility with our experiment environments, we replayed a 6-minute length trace segment and re-scaled its arrival intensity, as previous works have done [73, 24, 6, 45]. In the experiment, each inference request consisted of an initial prompt length of 64 tokens, generating 32 new tokens during the inference phase. The inference SLO required that the per-iteration latency did not exceed 50ms. We used a mini-batch size of 1024 tokens for the PEFT finetuning request. We conducted our evaluation using the LLaMA-30B model over 4 A100 GPUs.
In our case study, as shown in Figure 16, we observed that the arrival rate of the inference requests initially increased to a peak level after around two minutes, and then gradually decreased. FlexLLM could automatically detect the fluctuations in the workload and improve the ratio of inference tokens (vs fine-tuning tokens) in each iteration’s batch. This significantly increased inference throughput (from 3.5K to 4.7K) and GPU compute utilization (from 21.5% to 25%). The impact on the finetuning throughput was only minimal (less than 0.3K), which further highlights the potential waste of resources in a system that is solely focused on inference.
@@ 9 Related Work
Numerous ML serving systems [24, 25, 33, 51, 15] have been proposed in recent yearsto address various challenges such as latency prediction, system scalability, model performance, swapping [8], preemption [13], workload estimation [73], cost-effectiveness [71], and more. These systems include Nexus [56], which focuses on batching and scheduling multiple small DNN models within a GPU for higher execution throughput, and Gpulet [12], which improves resource assignment and utilization by creating multiple virtual GPUs within a single physical GPU. However, these systems are mainly designed for small CNN models. AlpaServe [37] extends the temporal sharing approach to large models via model parallelism across GPUs. Some concurrent PEFT serving systems like Punica [10] and S-LoRA [57] follow the design of PetS [77] and further improve it by leveraging new GPU kernels for latest PEFT methods like LoRA [29]. However, most of these existing works ignore the resource utilization issue in LLM inference systems. To the best of our knowledge, FlexLLM is the first PEFT co-serving system that concurrently executes finetuning and inference tasks and improves resource utilization without affecting inference latency.

### v2

@@ 6.2 Hybrid Token Scheduler
FlexLLM’s hybrid token scheduler coordinates token-level GPU execution in two stages. First, FlexLLM determines the scheduling of available inference requests based on specific scheduling policies. By default, FlexLLM adopts Orca’s iteration-level scheduling [84], which maintains a fixed maximum batch size and dynamically replaces each completed request with a new one whenever available. To further mitigate blocking caused by long input sequences, FlexLLM incorporates the chunked-prefill optimization [6].
Second, after determining the inference schedule for a given iteration, FlexLLM opportunistically appends as many finetuning tokens as possible—determined by the sliding window size $s$—to maximize GPU utilization. The number of finetuning tokens added is determined automatically using the formula $s=\operatornamewithlimits{arg\,max}f(c,s)\leq$SLO, where $f(\cdot,\cdot)$ is the latency estimation function and $c$ is the number of inference tokens scheduled in the current iteration. Here $f(\cdot,\cdot)$ is derived via offline profiling of the LLM’s execution [55]. Such window-based, best-effort scheduling mechanism ensures that FlexLLM can automatically adapt to inference workload fluctuations, maintaining inference SLOs while opportunistically allocating compute to finetuning tasks (§8.3).
Figure 9 illustrates the execution timeline of FlexLLM and shows an example a token scheduling plan for a finetuning mini-batch $A$, under a representative inference requests arrival pattern (i.e., $r_{0}$ to $r_{4}$). To preserve the semantics of finetuning, FlexLLM enforces the execution dependencies between the forward and backward passes of each mini-batch. During the forward pass, FlexLLM leverages fused GPU kernels to jointly process both inference and finetuning tokens. This approach avoids additional kernel launch overhead and is enabled by the shared token-wise computation logic between inference and finetuning. In particular, finetuning tokens follow the same execution pattern and causal masking rules as inference tokens during the prefill phase, guaranteeing compatibility at the kernel level. For the backward pass, FlexLLM adapts a space-sharing approach, using two separate GPU streams: one for inference tokens and another for finetuning tokens. This design allows concurrent execution without interference.
Figure 11: End-to-end comparison between co-serving
@@ 7 Implementation
FlexLLM is built on top of FlexFlow Serve [53] with 9K lines of C++ and CUDA code. It comes with both C++ and Python bindings, and its Python interface is fully compatible with the HuggingFace’s PEFT [50] and transformers [79] libraries. This means that users can easily serve models imported directly from HuggingFace, and even upload finetuned PEFTs to the HuggingFace Hub. Moreover, FlexLLM supports any PEFT models with LLAMA [73], GPT [12], OPT [90], Falcon [9], or MPT [70] backbones. In addition, we provide instructions on how to add support for custom models from HuggingFace.
@@
To support token-level finetuning, it is necessary to accumulate key-value gradients for every preceding token in a given sequence during backpropagation. FlexLLM utilizes static allocation to reserve space for key-value gradient accumulation. During backpropagation, the gradients for keys and values of a specific segment are obtained, and the position of this segment in the overall sequence is determined. FlexLLM then accumulates the new key-value gradients of all subsequently scheduled segments. When a segment is scheduled during backpropagation, it includes accumulated gradients from future tokens.
@@ 8 Evaluation
We evaluated the performance of FlexLLM by comparing it with existing approaches across three publicly available LLMs: LLaMA-3.1-8B [74], Qwen-2.5-14B [60], and Qwen-2.5-32B [60]. Pre-trained model weights were obtained from their respective HuggingFace repositories [1]. Since existing LLM serving systems only support base models and LoRA-based PEFT, our evaluation focuses on these two configurations to ensure fair comparison. Following prior work, we applied LoRA with a rank of 16 to the down projection layers in each multi-layer perceptron (MLP) block of the models. This setup yields 9.4M, 14.5M, and 25.16M trainable parameters for the three LLMs.
@@
Workload.
To synthesize the inference workloads, we sampled the prompt lengths and generation lengths of each request from the ShareGPT [72] dataset, following prior work [39, 95, 7]. However, ShareGPT does not include request arrival times. To address this limitation, we used arrival time data from a recent production trace [76] collected from ChatGPT serving on Azure. The original trace spans multiple days. To evaluate the online serving capabilities of FlexLLM, we sampled 20-minute intervals from the trace and replayed the requests, adjusting the arrival times to simulate different average arrival rates. We used a per-token inference latency of 50ms, 75ms, and 75ms as the SLO for inference requests, following prior work [7]. We set the time-to-first-token (TTFT) SLO to 5s per request to prevent excessive queueing.
We sampled each finetuning request from the NovaSky-AI reasoning dataset [71], and truncate each sequence length to a maximum of up to 8192 tokens. We used the Adam optimizer for training [38].
@@ 8.1 End-to-end Comparison
In this section, we compare FlexLLM’s co-serving approach with existing approaches that employ separate clusters for inference and finetuning tasks. In the separate cluster approach, we use vLLM [39] and LlamaFactory [94] as the inference and finetuning systems, respectively, as they represent state-of-the-art solutions for these workloads. In this experiment, we allocated 4, 8, and 16 A100 GPUs for LLaMA-3.1-8B, Qwen-2.5-14B and Qwen-2.5-32B, respectively, to serve inference and finetuning workloads.
Figure 10 shows the results. The separate cluster approach applied a tensor parallel degree of 1, 2, and 4 for the three models, respectively, creating an overall of four inference and finetuning pipelines. We explored various strategies to distribute resources between inference and finetuning tasks. For example, a configuration of 25% vLLM - 75% LlamaFactory indicates that 25% of the resources (i.e., one pipeline) are dedicated to inference, while 75% (i.e., 3 pipelines) are for finetuning.
Across all three models, configurations using one or two pipelines for inference—25% vLLM and 50% vLLM—are only capable of managing lightweight inference workloads, achieving limited inference throughput and SLO attainment under more intensive workloads. Conversely, dedicating three pipelines to inference (75% vLLM - 25% LlamaFactory) maintains high SLO attainment for inference tasks but results in suboptimal finetuning throughput.
FlexLLM’s co-serving approach maintains near-optimal SLO attainment but without compromising finetuning performance. By prioritizing latency-sensitive inference tasks in each co-serving iteration and opportunistically appending as many finetuning tokens as possible, FlexLLM maximizes GPU utilization and improves finetuning throughput.
Across all three models, FlexLLM matches the 75% vLLM - 25% LlamaFactory configuration in inference SLO attainment (at or above 90% even at 20 req/s) and inference throughput, while dramatically improving finetuning throughput. Specifically, under heavy inference loads (20 req/s), FlexLLM sustains finetuning throughputs of 7.2K, 2.2K and 2.2K tokens/s for LLaMA-3.1-8B, Qwen-2.5-14B, and Qwen-2.5-32B respectively, compared to only 3.8K, 1.0K, and 0.5K tokens/s in the 75% vLLM - 25% LlamaFactory setup—i.e. a 1.9$\times$-4.8$\times$ improvement. Under light inference loads (4.0 req/s), FlexLLM sustains finetuning throughputs of 9.4K, 3.7K, and 3.2K tokens/s for the same models, translating to a 2.5$\times$-6.8$\times$ improvement over the 75% vLLM - 25% LlamaFactory setup.
8.2 GPU Scheduling
This section compares FlexLLM’s co-serving mechanism with two commonly used GPU scheduling strategies: temporal sharing and spatial sharing. Temporal sharing interleaves the execution of inference and finetuning tasks over time, as illustrated in Figure 1. Spatial sharing simultaneously launches inference and finetuning computations using separate CUDA resources. Both strategies are implemented on top of FlexLLM by replacing the co-serving mechanism with their respective scheduling strategies. To ensure a fair comparison, all memory optimizations are enabled in each baseline.
Figure 11 shows the results. For temporal sharing, interleaving one inference iteration with one finetuning iteration would violate the SLOs for nearly all inference requests. This is because each inference iteration is expected to complete within tens of milliseconds, whereas a finetuning iteration requires several seconds. To mitigate this issue, in the temporal sharing experiments, we interleave each finetuning iteration with $n$ inference iterations, where $n$ is the inference frequency shown in the figure.
While adopting a low inference frequency in temporal sharing (i.e., frequency = 64) maximizes finetuning throughput, it adversely impacts the SLO attainment and throughput for inference requests. Employing a frequency of 128 enables temporal sharing to match the inference throughput and SLO attainment rate of co-serving. However, it also reduces finetuning throughput by 1.2-1.8$\times$ compared to co-serving.
Temporal sharing allocates separate GPU resources for inference and finetuning, achieving comparable finetuning throughput to co-serving. However, it remains suboptimal in SLO attainment under heavy inference workloads due to the interference between inference and finetuning tasks.
(a) Arrival Inference Requests
(b) System throughput
Figure 12: Case study of FlexLLM’s system throughput for fluctuating inference workload.
8.3 Case Study
This section describes a case study of FlexLLM using a a 10-min interval of the BurstGPT trace with Qwen-2.5-14B model. To ensure compatibility with our experiment environments, we replayed this trace segment and re-scaled its arrival intensity, as previous works have done [89, 29, 8, 54]. In the experiment, each inference request and inference request are sampled from the ShareGPT dataset and the Sky-T1 dataset respectively.
In our case study, as shown in Figure 12, we observed that the arrival rate of the inference requests initially increased to a peak level after around 90 seconds, and then gradually decreased with some peaks. FlexLLM could automatically detect the fluctuations in the workload and improve the ratio of inference tokens (vs fine-tuning tokens) in each iteration’s batch. This significantly increased inference throughput from a few hundreds to 2.25K.
@@ 8.4 Memory Optimization
Figure 13: Ablation study of memory optimizations.
FlexLLM introduces a series of memory optimizations to reduce the memory overhead of PEFT finetuning and allow co-serving finetuning with inference workloads. To understand the effectiveness of FlexLLM memory optimizations, we perform an ablation study that incrementally turns off these optimizations (graph pruning, rematerialization, and token-level finetuning) and measures the memory requirements. Figure 13 shows the activation memory requirements of FlexLLM for different finetuning methods on a 70B LLM and a sequence length of 1024. FlexLLM saves $85\%$-$87\%$ memory requirements for activations compared to existing approaches. The major improvement comes from graph pruning. With graph pruning only, FlexLLM achieves $71\%$-$74\%$ activation memory overhead reduction. Rematerialization and token-level finetuning further reduce memory overhead by 00-$8\%$ and $4\%$-$10\%$, respectively.
@@ 9 Related Work
Numerous ML serving systems [29, 30, 40, 62, 18] have been proposed in recent years to address various challenges such as latency prediction, system scalability, model performance, swapping [10], preemption [16], workload estimation [89], cost-effectiveness [87], and more. These systems include Nexus [64], which focuses on batching and scheduling multiple small DNN models within a GPU for higher execution throughput, and Gpulet [15], which improves resource assignment and utilization by creating multiple virtual GPUs within a single physical GPU. However, these systems are mainly designed for small CNN models. Recently, LLM serving systems attracts lots of attention [52], such as vLLM [39], Sarathi [7], and SpecInfer [53]. Among them, AlpaServe [45] extends the temporal sharing approach to LLMs via model parallelism across GPUs. Some concurrent PEFT serving systems like Punica [13] and S-LoRA [66] follow the design of PetS [97] and further improve it by leveraging new GPU kernels for latest PEFT methods like LoRA [35]. However, most of these existing works ignore the resource utilization issue in LLM inference systems. To the best of our knowledge, FlexLLM is the first PEFT co-serving system that concurrently executes finetuning and inference tasks and improves resource utilization without affecting inference latency.